A Complete Bibliography of Publications in Biometrika for the decade 1940–1949

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.03

Title word cross-reference

(n × 2) [Hal40a, Hal45c]. 2 × 2 [Bar47c, Bar47a, Fin48, Moo49, Pat48, Pat59, Pea47, PM48, Sil49]. b_{2} [GW47, GW50, HL40]. χ [Dav48]. χ^{2} [Dav47a, Hal40a, Hal40b, Hal45c, Irw49, Lan49b, Lan50b, Mer41, Sea48, Tho41, Tho47, Wis47, Ano47c, Hal45a, Pat49]. F [Dav49c, Pat49]. k [Dav49d, Sea45]. n [PH42]. p [Sch44]. r [Hal45a]. s [Bat49]. \sqrt{b_{1}} [Gea47a]. t [Ano47c, Bal46, Gay49, Ghu49, JW40, Lor47, Lor52, Mer42, Wis47]. τ [Hof47, Ken47, Sil47]. Z [Ano47c, Wis47, Aro47, Car47, Dav49c].

-Distribution [Aro47, Bal46, Car47, JW40, Mer42, Pat49]. -Fold [Hal40a, Hal45c]. -Statistics [Dav49d]. -Test [Ghu49, Lor52, Lor47].

1939 [M.41].

Absence [Hal45a]. Accuracy [AW49]. Admitting [Huz49]. Advanced [W.48, Wel45]. Advances [H.47]. Age [GM40]. Air [Mor48c]. Alternative [Joh48a]. Alternatives [Bat48, Dav47d, Moo49]. Analysis [Bea42, FS48, Har49, Hsu40, Hsu41a, Joh48a, Nai48b, Nic41, Rao48, Ste48, Toc49, Wil40, Yat48a, Yat48b, J.48]. Annual [Bat41, Sea45]. any [Gay49]. Apodemus [HP44, HP46]. Appendix [Har45]. Applicable [Bea42]. Application [Dav49d, Gar41, GW50, Iye49, PW40, Sic49, GW47]. Applications [JW40, Pat49]. Applied [Cha46]. Approach [Kem47]. Approximate [Tho47]. Approximation [Car47, Har48a]. Approximations [Mer41, Sil49]. Argument [Ney41]. Arising [Arm49, Asp48, Iye49, Lan49a, Lan50a]. Arrangements [BKL48]. Association [Hal45a]. Asymmetrical [Ghu49]. Asymptotic [Cox48]. Asymptotical [Elf47]. Autocorrelations [Que47]. Autoregressive [Ken44, Ken49c, Que47].

Back [Ano40a, Ano41a, Ano41b, Ano42a, Ano43a, Ano44a, Ano45a, Ano46a, Ano47a, Ano47b, Ano48a, Ano48b, Ano49a, Ano49b]. Based [Joh49a, Yat48a, Yat48b]. Becomes [Bea42]. Being [Ano41f]. Below [Sea45]. Beta [MT43, Sch44, TPCH41, Tho47]. Beta-Function [TPCH41]. Between [BL48, Dan44, Dav48, Fin47, Nic41, OJ48, Pat48, Pat59, Sim42, Whi47]. Bibliography [M.41]. Biennial [Bat41]. Binomial [Ans48, Hal40a, Sch44, Tho49]. Birth [Ken48b]. Bivariate [Joh49a]. Bombing [Gar47]. Book [Dav40, Dav47b, Dav47c, Dav49b, Dav49a, H.47, J.48, M.41, P.42, Smi49, W.48, Wel45, Wil41, Yul49]. Books [Ano41c, Ano43b, Ano44b, P.42]. Both [Ken47]. Breadth [Sch40]. British [Mor48c]. Buros [P.42].

C [Dav47b, SBL79]. Calculate [Sch44]. Calculation [FS48, GW47, GW50, Har45, Que47]. Canonical [Bar41, Hsu41b]. Carota [Bat41]. Carrot [Bat41]. Castle [GM40, Lit43]. Cells [Dav48]. Central [Bat49, JW40, Pat49]. Certain [CH41, Lan49b, Lan50b, Les45, Sch40]. Chance [Ano41f]. Characteristic [Bat49]. Characters [Yat48a, Yat48b]. Chinese [Woo43]. Choice [Pea47]. Civilians [Mor48c]. Class [Box49, Whi49]. Classed [Pea47]. Closeness [Gea44]. Clover [PW40]. Coefficient [Hof47, Mor48b, Sil47]. Coefficients [CH41, Que48]. College [Ano41f]. Combination [Lan49a, Lan50a]. Comments [SBL79]. Comparative [Sil49]. Comparing [Asp48]. Comparison [Arm47, Bat41, Car49, Gea44]. Comparisons [AW49, KS40, Mor47a, Mor48c]. Completely [BL48]. Components [Irw49]. Composed [Sea45]. Computation [GW47, GW50]. Conceptions [Ano41f]. Concepts [Gea44]. Concluded [Gre43]. Conduct [Ano41f]. Confidence [Ney41, Sch44]. Confounding [Kem47]. Connected [Ken41a]. Consistent [Gea44]. Contain [Ken47]. Containing [Sil47]. Contingency [Fin48, Yat48a, Yat48b]. Continuation [Ken48c]. Continuous [Jon48].

Control [Tip44]. Correlation [Dan44, DK47, Hal49, Hof47, Ken42, Ken49b, Mor48a, Mor48b, OJ48, Que48, Sil47, Whi47, Whi49, Dav40]. Correlations [Bar41, DK47, Dan48, Dav48, Hsu41b]. Corrigenda [Ano47c, Ano48c, Ano48d, GW50, Lan50a, Lan50b, Lor52, Pat59, Yat48b]. Counting [Arm49]. Courbes [Dum48]. Course [Dav47b]. Cramér [Dav47c]. Criteria [Box49]. Criterion [Wil40]. Cubes [BL48]. Cumulants [Ano47c, Aro47, Hal40a, Hal41, Hal42a, Wis49, Wis47]. Curves [Gar41, Joh49b, Ken41b, Dum48].

D [J.48]. Daily [Rus45]. Data [Ans48, Bea42, Lan49a, Lan50a, Pea47, Toc49]. Daucus [Bat41]. David [Smi49]. Definitions [Ano41f]. Delivered [Ano41f]. Depth [Sch40]. Derivation [HL40, Lan49b, Lan50b]. Derived [GW47, GW50]. Design [PB46, Pla46]. Development [Asp48, Ken48b]. Deviate [Nai48a]. Deviation [Ano45c, Ano48c, Ano48d, Gea36, God45, God48, Har44, Har45, Lor52, Nai47, Nai49, Pla47b, SBL79, Tru40, Lor47]. Dichotomous [Whi47]. Difference [Joh48b, Pat48, Pat59]. Different [Wel47]. Discrete [Lan49a, Lan49b, Lan50a, Lan50b, Mor49]. Dispersion [God49, Jon46]. Distribution [Ait49, Aro47, Bal46, Box49, Car47, Cox48, Dan41, Elf47, Gay49, Gea47a, God45, Hal40a, Hal41, Hal42a, Hal48a, Har44, Har45, Hof47, HL40, Hsu41b, JW40, Ken48a, Mer41, Mer42, MT43, Nai48a, Nic41, Pat49, She49, Sil47, Tho41, Tho47, Tru40, Wis48a, Wis48b]. Distributions [Ano47c, Dav49c, Fin49a, Gea43, Hal42b, Har48a, Huz49, Iye49, Joh49a, Lan49a, Lan49b, Lan50a, Lan50b, Les42, Wis49, Wis47]. Dodder [PW40]. Dorset [GM40]. Dosage [Gar41]. Dosage-Mortality [Gar41]. Dose [Fin47]. Dr [Ken48c]. Drawn [Gay49, GW47, Gea47a, GW50]. Duck [Ren43]. Dynasty [BM47].

E. [Bar47a]. Each [Sea45]. Early [Yul49]. Ecology [Tho49]. Editorial [Pea45]. Efficiency [Gea44, She49]. Eggs [Ren43]. Egyptian [BM47]. Eleventh [BM47]. Elimination [Har44, Que49]. Entomological [Bea42]. Entrants [Sea45]. Equality [Pla47a]. Equations [Lin46]. Errors [Har48a, Rus45]. Estimate [God45, Har45]. Estimated [AW49, DJ48, TM46]. Estimates [Gea44]. Estimating [Fin49b, Hal45b]. Estimation [Ans49, Car49, Fin47, Fin49a, God49, Har48b, Jon46]. Events [Pea48]. Exact [Pla47a]. Examination [Asp48]. Exists [DK47]. Expectations [Hal40b, Hal45c]. Experiment [PM48, Ste48]. Experiments [Bea42, BKL48, Kem47, PB46]. Extreme [Nai48a, Nai48b]. Extremes [Dan41].

F [MT43, Smi49]. Facial [Sch40]. Factorial [BKL48, Kem47, Ste48]. Factors [BKL48]. Farr [Gre41, Gre42, Gre43]. Fiducial [Ney41]. Field [Bea42, HP44, HP46]. Fifth [HL40]. Figures [Gar47]. Finite [Arm47, BL48]. Finney [J.48]. First [Ano41f, Dav47b, Iye49, BM47]. Fisher [Aro47, Dav49d, Fin48, Ken48a]. Fit [Dav47a]. Fold [Hal40a, Hal45c]. Force [Mor48c]. Form [Nai48a, Nai48b]. Formula [Asp48, Hun40]. Formulae [Tho47]. Fourfold [Hal45a]. Fowler [Gre48]. Fractional [BKL48, Kem47]. French [Dum48]. Frequence [Dum48]. Frequencies [Fin49b, Hal45b, Hal48b]. Frequency [Gea47a, Joh49b, Les42, Nic41, Sic49, Dum48]. Frequency-Moments [Sic49]. Fresh [Sim42]. Front [Ano40b, Ano41d, Ano41e, Ano42b, Ano43c, Ano44c, Ano45b, Ano46b, Ano47d, Ano47e, Ano48e, Ano48f, Ano49c, Ano49d]. Function [Bat48, Bat49, Dav47d, Hsu41a, Pat48, Pat59, PM48, Sch44, Sil49, Sim42, TPCH41, Tho47]. Functions [DK49]. Fundamental [Ano41f]. Fundamentals [Dav49a]. Further [Asp48, God48, Les48, Nai49, Pea41].

G [W.48, Wel45]. Geary [SBL79]. General [Box49]. Generalization [Ken41a, Tho49, Wel47]. Generalizations [Pla46]. Generalized [Hsu40]. Generated [Joh49b]. Generation [Ken48b]. Genetics [H.47]. Geometrical [Fog48, Gar47, Nic41]. German [Mor48c]. Given [Gum42, Hal42a]. Goodness [Dav47a]. Graunt [Gre41, Gre42, Gre43]. Gresham [Ano41f]. Grouped [Toc49]. Groupings [Yat48a, Yat48b]. Groups [BL48]. Growth [HP44, HP46, Ken48a].

H [Dav47c]. Haldane [Pea48]. Harmonic [CH41, Har49]. Harold [Dav49b, Wil41]. Hatched [Ren43]. Heights [Mor48c]. Hill [Lit43]. Historical [Pla49]. Homogeneity [Hal40b, Hal45c, Har40, PW40, TM46]. Human [GM40, Lit43, M.41]. Humerus [Woo43]. Hyper [BL48]. Hyper-Cubes [BL48]. Hypotheses [Gum42, Pea42].

II [Mor48b]. Illustrated [Pea47]. Important [Sim41]. Incomplete [Sch44, TPCH41, Tho47]. Independent [Har48a, Hof47]. Individual [Fin47]. Industrial [Tip44]. Inequalities [Les42, Win46]. Infectiousness [Gre49]. Inflorescences [Bat41]. Information [Ano40c, Ano42c, Ano46c, Ano47f, Ano48g, Ano49e]. Integral [Ano45c, DJ48, Nic43, PH42, PH43]. Internal [Har48b]. Interpolation [CH41, Sim42]. Interpretation [Bar40, Pea47]. Intervals [Ney41, Sch44]. Intra [Whi49]. Intra-Class [Whi49]. Introductory [Ano41f]. Invariants [Ken41b]. Inverted [MT43]. Involved [Wel47]. Involves [AW49]. Involving [Moo49]. Iron [GM40].

J [J.48]. James [Gre48]. Jeffreys [Dav49b, Wil41]. Jones [Ken48c].

K. [Dum48]. Karl [Yul49, Ano41f, Yul49]. Kelley [Dav49a]. Kendall [Sil47, W.48, Wel45]. Krisen [P.42]. Krogman [M.41].

L [Dav49a]. Lagrangian [CH41]. Large [Ans49, Jon46]. Large-Sample [Ans49]. Lattice [Iye49]. Law [Wis48b, Wis48a]. Laws [Ano41f]. Leading [Ken48a]. Least [Har48b, Lin46, Pla49]. Lectures [Ano41f]. Length [Wil40]. Level [Bar47b, Sea45]. Levels [BKL48, Sim42]. Likelihood [Box49, Gar41]. Limit [Tho49]. Limiting [Hsu41b]. Limits [Pla47b]. Linear [Bat49, God49, Har48b, Que47]. Literary [Wil40]. Logarithmic [Ano47c, Ken48a, Wis47]. Long [HP44, HP46]. Long-Tailed [HP44, HP46]. Longest [Bat48]. Lowest [Sea45].

M [M.41, W.48]. Maiden [GM40]. Man [Sch40]. Mathematical [Dav40, Dav47b, Dav47c]. Mathematics [Les45, Les48, Sea45]. Matrices [Les45, Les48]. Matter [Ano40a, Ano40b, Ano41a, Ano41b, Ano41d, Ano41e, Ano42a, Ano42b, Ano43a, Ano43c, Ano44a, Ano44c, Ano45a, Ano45b, Ano46a, Ano46b, Ano47a, Ano47b, Ano47d, Ano47e, Ano48a, Ano48b, Ano48e, Ano48f, Ano49a, Ano49b, Ano49c, Ano49d]. Maurice [Wel45]. Maximum [Gar41]. Mean [Ano45c, Ano48c, Ano48d, Asp48, Gea36, God45, God48, Hal40b, Har45, Nai47, Nai48a, Nai48b, Nai49, Pla47b, SBL79, Tru40, Win46]. Meaning [Bar47b]. Measles [Gre49]. Measurement [Rus45]. Measurements [Sch40]. Measures [Dan44]. Median [Hal42a, Nai47, Nai49, Hal48a]. Medical [Gre41, Gre42, Gre43]. Method [Fin49b, Fog48, Hal45b, Joh48b, Jon46, KS40, Mor47a, Pla49, Que49, She49, Sic49]. Methodology [P.42]. Methods [Dav47c, Joh49b, Smi49]. Minimum [Gea43]. Mode [Hal42a]. Modes [Ken48a]. Moment [GW47, GW50, Ken49b, Mor48a, Wis48a, Wis48b]. Moments [Ano48c, Dav49c, Gea36, GW47, GW50, Hal40a, Hal42b, Hal45a, HK47, HL40, Iye49, She49, Sic49]. Monozygotic [Coc40]. Morphology [M.41]. Mortality [Gar41]. Mouse [HP44, HP46]. Multifactorial [PB46, Pla46]. Multinomial [Wis49]. Multivariate [Hal48a, Les42, Rao48, Wis49].

N [Smi49]. Nearly [Hal42a]. Negative [Ans48]. Negative-Binomial [Ans48]. Neyman [She49]. Non [Bat49, Gay49, Hal49, Har48b, JW40, Pat49, Ste48]. Non-Central [Bat49, JW40, Pat49]. Non-Linear [Har48b]. Non-Normal [Gay49, Hal49]. Non-Orthogonal [Ste48]. Normal [Ano45c, Ano48c, Bat49, Elf47, Gay49, Gea36, Gea43, GW47, Gea47a, GW50, God45, Hal42b, Hal42a, Hal49, Har45, HL40, PH42]. Normality [GW47, Gea47b, GW50]. Note [Aro47, Bar47a, Bar40, Cox48, Dav49d, God48, Hal48a, Hal49, Har45, Irw49, Nai47, Nai49, Pea41, Pea45, Pea48, Pla49, Sch44, Sea48, Sil49, Toc49, Wil40, Wis48a]. Notes [Les48, Pea42, Que47]. Number [Sea45]. Numerical [HK47, Mer41].

Observations [Car49, PH42]. Observed [Hal48b]. Obtained [God45]. One [Whi47]. Optimum [PB46, Sim41]. Order [Wis48a, Wis48b]. Orthogonal [BL48, Ste48]. Oscar [P.42]. Other [Coc40, GM40, Mor48c, Whi47]. Overlap [Arm49, Gar47].

Paired [KS40, Mor47a]. Paper [Bar47a, Ken48c, Pea48, SBL79]. Papers [Yul49]. Parameter [Gea44, Jon48]. Parameters [DJ48, Fin49a, Har48b]. Parent [Har44]. Parental [DK47]. Part [DK49]. Partial [Ken42]. Particle [Arm49]. Partition [Lan49b, Lan50b]. Pearson [Yul49, Ano41f, Bar47a, Dum48, Ken41b, Yul49]. Percentage [Bal46, Car47, CH41, Mer41, Mer42, MT43, Sch44, TPCH41, Tho41, Tho47]. Periods [GM40]. Permutations [Dan44]. Place [Lor52, Lor47]. Points [Bal46, Car47, CH41, Iye49, Mer41, Mer42, MT43, Sch44, TPCH41, Tho41, Tho47]. Poisson [Ans48, PW40, Tho49]. Population [Arm47, Elf47, Gea47a, Ghu49, God45, Har44, HL40, Ken48a, Les45, Les48, PH42, Sea45, Vaj47, Wel47]. Populations [Jon48, Sic49]. Power [Bat48, Dav47d, Hsu41a, Pat48, Pat59, Sil49, PM48]. Powers [Hal42b]. Precision [Hal48b]. Principles [Dav40]. Probabilities [Lan49a, Lan50a]. Probability [Ano45c, Ano48d, DJ48, Gum49, Iye49, Ken49a, Nic43, PH42, PH43, Sim42, Dav49b, Wil41, Smi49]. Probit [FS48, Toc49, J.48]. Probits [FS48]. Problem [Arm49, Asp48, Gar47, HK47, Wel47]. Problems [Ken45]. Process [Ken48b]. Processes [Mor49, Tip44]. Product [Ken49b, Mor48a]. Product-Moment [Ken49b, Mor48a]. Products [Hal42b]. Professor [Pea48]. Projective [Sch40]. Proof [Ken41a]. Proofs [Wis48a, Wis48b]. Properties [Pea41]. Property [Dan41, Dan48, Huz49, Sim41]. Proportions [Pat48, Pat59]. Psychology [Cha46]. Puerperal [Rus45].

Quality [Tip44]. Quantal [Fin47]. Quantitative [Yat48a, Yat48b]. Quasi [Bar40, Gea43]. Quasi-Normal [Gea43]. Quasi-Sufficiency [Bar40].

R. [Ken48a, SBL79]. Random [Arm47, Gay49, GW47, Gea47a, GW50, Har42, Har44]. Randomized [PM48]. Randomness [Bat48, Dav47d, Ken41c, Moo49]. Range [Cox48, Elf47, Gea43, Gum49, Har42, Lor52, PH42, PH43, Pla47b, Win46, Lor47]. Rank [DK47, Dan48, Hof47, Ken42, Ken49b, Mor48a, Sil47, Whi47, Whi49]. Ranked [Whi47]. Ranking [Ken45]. Rankings [Ken47, Sil47]. Rare [Pea48]. Ratio [Ano48c, Gea36, Nic41, Pla47b]. Received [Ano41c, Ano43b, Ano44b]. Reconciliation [Ken49a]. Recorded [Mor48c]. Records [Fin47]. Recruited [Sea45]. Recurrence [Ken41b]. Reference [Gar47]. Regarding [Pea48]. Regressions [Car49]. Related [Car49]. Relation [Ano41f, Dan44, Ken41b]. Relations [Ken41a]. Relationship [BL48, Fin47]. Remainder [Hun40]. Remains [GM40]. Replication [BKL48, Kem47]. Research [P.42]. Residual [Car49]. Response [Fin47]. Restraints [Bat49]. Results [PW40, Que48]. Reversals [Coc40]. Review [Ano44b, Dav40, Dav47b, Dav47c, Dav49b, Dav49a, H.47, J.48, M.41, P.42, Smi49, W.48, Wel45, Wil41, Yul49]. Reviews [P.42]. Role [Ken48b]. Routine [Jon46, Rus45]. Royal [Mor48c]. Run [Bat48].

S [Bar47a]. Sakkara [BM47]. Sample [Ans49, Dan44, DJ48, Har44, Nai48a, Sch44]. Sample-Distribution [Har44]. Samples [Ano45c, Ano48c, Elf47, Gay49, Gea36, GW47, Gea47a, GW50, God45, Har42, Har45, HL40, Jon46, PH42, PW40, Tru40]. Sampling [Arm47, Fog48, Jon48]. Scarborough [Lit43]. Schemes [Que47]. Second [Iye49, P.42, Wis48a, Wis48b]. Seed [PW40]. Semi [Ken41b, Vaj47]. Semi-Invariants [Ken41b]. Semi-Stationary [Vaj47]. Sentence [Wil40]. Sentence-Length [Wil40]. Separately [AW49]. Sequence [Bat48, Dav47d, Moo49]. Sequential [Ans49]. Serial [Mor48b, Que48]. Series [Ano41f, BM47, Ken41a, Ken44, Ken48a, Ken49c, Lit43, Mor47b, Mor48b, Mor48c, OJ48, PW40]. Set [Har40, TM46]. sets [Car49]. Seventh [GW47, GW50]. Several [Wel47]. Significance [Bar47b, Bar47c, Bar41, DK47, Fin48, Har49, Joh48b, Mor48b, OJ48, Rao48]. Simple [Gum42, Joh49a, Kem47]. Simultaneous [Tru40]. Sixth [HL40]. Size [Gay49]. Sizes [Gea47a]. Skeleton [Sch40]. Skulls [BM47, Lit43]. Small [Hal40b, Hal45c, Hal48b, Sch44]. Smooth [Dav47a, Sea48]. Solution [HK47, Lin46]. Some [Ken48a, Les48, Lin46, Mor47b, Mor48b, Pla46, Que48, Iye49]. Spectral [Mor49]. Square [Hal41, Har48b, Nai48b]. Squares [Bat49, BL48, Lin46, Pla49]. Standard [Ano48c, Gea36, Har44, Lor52, Pla47b, Sim42, Tru40, Lor47]. Standardized [SBL79]. Standpoint [Hsu41a]. Stationary [Sea45, Vaj47]. Statistic [Wis48a]. Statistical [Bar41, Cha46, GW47, GW50, P.42, Pea41, Pea42, Pea47, Sim41, Smi49, Ste48, Wil40, Yul49]. Statistical-Techniques [Cha46]. Statistics [Ait49, Dav49d, God49, Gre41, Gre42, Gre43, Har44, Huz49, Wis48b, Dav47b, Dav47c, Dav49a, W.48, Wel45]. Stirling [Hun40]. Stochastic [Ken48b, Mor49]. Strata [Sea45]. Stratified [Arm47, Vaj47]. Stratum [Sea45]. Student [Gay49, Ghu49, Wel47]. Studentization [Har44]. Studentized [Nai48a, Nai48b, PH43]. Study [BM47, Lit43, Woo43]. Style [Wil40]. Subdividsion [Irw49]. Subject [Bat49, Tip44]. Sufficiency [Bar40]. Sufficient [Huz49]. Sum [Bat49]. Supported [Sea45]. Survival [HP44, HP46]. Sylvaticus [HP44, HP46]. Symmetric [DK49]. Symmetry [Coc40]. Systematic [God49, Jon48]. Systems [Joh48a, Joh49a, Joh49b].

T [Dav49a]. Table [Bal46, CH41, FS48, Hal40a, Hal45a, Hal45c, Mer42, Moo49, Pat48, Pat59, Pea47, Sim42, Tho41]. Tables [Ano45c, AW49, Bar47c, Bar47a, CH41, DK49, Fin48, Gum49, Ken49c, MT43, PH43, PM48, Sch44, TPCH41, TM46, Yat48a, Yat48b]. Tailed [HP44, HP46]. Techniques [Cha46]. Terms [Win46]. Test [Bat48, Dav47a, Fin48, Ghu49, Hal40b, Hal45c, Lor52, Moo49, Nai48b, Pat48, Pat59, PM48, Pla47a, Sea48, Lor47]. Testing [Gea47b, Har40, OJ48, Pea42, PW40, Que48, TM46]. Tests [Bar47c, Dav47d, GW47, GW50, Gum42, Har49, Joh48b, Pea41, Pea47, Rao48, Sim41]. Tetrachoric [Ken41a]. Thebes [BM47]. their [Iye49, Pat49]. Theorems [Mor47b, Mor48b]. Theories [Ken49a]. Theory [Ans49, Box49, Dav40, Fog48, Ken41c, Mor49, Ney41, W.48, Wel45, Dav49b, Wil41]. there [Car49]. Thoery [Smi49]. Thought [Ano41f]. Tie [Sil47]. Ties [Ken45, Ken47]. Time [Ken44, Ken48b, Mor47b, Mor48b, OJ48]. Tocher [Gre48]. Tolerance [Fin49a]. Transformation [Ans48, Bea42, DJ48]. Translation [Joh49a, Joh49b]. Transposition [Coc40]. Treatment [Ken45, Pea48]. Trend [Que49]. Trends [Tip44]. Tri [Ste48]. Tri-Factorial [Ste48]. Trial [Sil49]. Twins [Coc40]. Two [Asp48, AW49, Moo49, Nic41, Nic43, Pat48, Pat59, Sim41, Whi47, BKL48, Car49]. Type [She49, Sic49].

Unhatched [Ren43]. Uniform [Sea45]. Universal [GW47, GW50]. Universe [Dan44, GW47, GW50]. Universes [Gay49]. Unrestricted [Arm47]. Use [AW49, Ghu49, Hal45c, Les45, Les48, Lor52, Sch44, Lor47, Tho49]. Used [Hal40b]. Useful [Jon46]. Uterus [Rus45].

Values [Asp48, Hal48b]. Variable [Ken48b]. Variables [Nic41, Nic43, Whi47]. Variance [Bea42, Gar47, Hal40b, Hsu40, Hsu41a, Joh48a, Ken47, Nai48b, Tru40]. Variances [AW49, Har40, Pla47a, TM46, Wel47]. Variate [Hal41, Joh48b]. Variates [Bat49, Hal42b, Har48a, Hof47]. Variation [HP44, HP46]. Variations [Ren43]. VII [Sic49]. Viscera [Coc40]. Volume [Ano40c, Ano42c, Ano46c, Ano47f, Ano48g, Ano49e].

Wandering [HP44, HP46]. Weatherburn [Dav47b]. Weighted [Bat49]. Weights [FS48, Mor48c, Ren43]. Where [DK47, Car49]. Which [Whi47]. Whose [AW49]. Wild [Bat41]. Wilton [M.41]. Wishart [Ait49]. Working [FS48].

Yates [Fin48]. Yearbook [P.42].

References

Aitken:1949:WDS

[Ait49] A. C. Aitken. On the Wishart distribution in statistics. Biometrika, 36(1/2):59–62, June 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

         stable/2332529.

Anonymous:1940:BM

[Ano40a] Anonymous. Back matter. Biometrika, 31(3/4):??, March 1940. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

         URL http://www.jstor.org/stable/2332621.

Anonymous:1940:FM

[Ano40b] Anonymous. Front matter. Biometrika, 31(3/4):??, March 1940. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332604.

Anonymous:1940:VI

[Ano40c] Anonymous. Volume information. Biometrika, 31(3/4):i–viii, March 1940. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332603.

Anonymous:1941:BMa

[Ano41a] Anonymous. Back matter. Biometrika, 32(1):??, January 1941. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332256.

Anonymous:1941:BMb

[Ano41b] Anonymous. Back matter. Biometrika, 32(2):??, October 1941. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332217.

Anonymous:1941:BR

[Ano41c] Anonymous. Books received. Biometrika, 32(1):88, January 1941. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332255.

Anonymous:1941:FMa

[Ano41d] Anonymous. Front matter. Biometrika, 32(1):??, January 1941. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332244.

Anonymous:1941:FMb

[Ano41e] Anonymous. Front matter. Biometrika, 32(2):??, October 1941. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332204.

Anonymous:1941:LCR

[Ano41f] Anonymous. The laws of chance, in relation to thought and conduct: Introductory, definitions and fundamental conceptions being: the first of a series of lectures delivered by Karl Pearson at Gresham College in 1892. Biometrika, 32(2):89–100, October 1941. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332205.

Anonymous:1942:BM

[Ano42a] Anonymous. Back matter. Biometrika, 32(3/4):??, April 1942. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332139.

Anonymous:1942:FM

[Ano42b] Anonymous. Front matter. Biometrika, 32(3/4):??, April 1942. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332125.

Anonymous:1942:VI

[Ano42c] Anonymous. Volume information. Biometrika, 32(3/4):i–vii, April 1942. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332124.

Anonymous:1943:BM

[Ano43a] Anonymous. Back matter. Biometrika, 33(1):??, April 1943. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333625.

Anonymous:1943:BR

[Ano43b] Anonymous. Books received. Biometrika, 33(1):104, April 1943. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333624.

Anonymous:1943:FM

[Ano43c] Anonymous. Front matter. Biometrika, 33(1):??, April 1943. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333615.

Anonymous:1944:BM

[Ano44a] Anonymous. Back matter. Biometrika, 33(2):??, August 1944. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334118.

Anonymous:1944:BRR

[Ano44b] Anonymous. Books received for review. Biometrika, 33(2):182, August 1944. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334117.

Anonymous:1944:FM

[Ano44c] Anonymous. Front matter. Biometrika, 33(2):??, August 1944. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334109.

Anonymous:1945:BM

[Ano45a] Anonymous. Back matter. Biometrika, 33(3):??, November 1945. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332309.

Anonymous:1945:FM

[Ano45b] Anonymous. Front matter. Biometrika, 33(3):??, November 1945. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332296.

Anonymous:1945:TPI

[Ano45c] Anonymous. Tables of the probability integral of the mean deviation in normal samples. Biometrika, 33(3):259–265, November 1945. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332307.

Anonymous:1946:BM

[Ano46a] Anonymous. Back matter. Biometrika, 33(4):??, June 1946. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332200.

Anonymous:1946:FM

[Ano46b] Anonymous. Front matter. Biometrika, 33(4):??, June 1946. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332190.

Anonymous:1946:VI

[Ano46c] Anonymous. Volume information. Biometrika, 33(4):i–vii, June 1946. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332189.

Anonymous:1947:BMa

[Ano47a] Anonymous. Back matter. Biometrika, 34(1/2):??, January 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332522.

Anonymous:1947:BMb

[Ano47b] Anonymous. Back matter. Biometrika, 34(3/4):??, December 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332456.

Anonymous:1947:CCL

[Ano47c] Anonymous. Corrigenda: “The Cumulants of the Z and of the Logarithmic χ^{2} and t Distributions”. Biometrika, 34(3/4):374, December 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332455. See

[Wis47].

Anonymous:1947:FMa

[Ano47d] Anonymous. Front matter. Biometrika, 34(1/2):??, January 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332507.

Anonymous:1947:FMb

[Ano47e] Anonymous. Front matter. Biometrika, 34(3/4):??, December 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332431.

Anonymous:1947:VI

[Ano47f] Anonymous. Volume information. Biometrika, 34(3/4):i–vi, December 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332430.

Anonymous:1948:BMa

[Ano48a] Anonymous. Back matter. Biometrika, 35(1/2):??, May 1948. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332644.

Anonymous:1948:BMb

[Ano48b] Anonymous. Back matter. Biometrika, 35(3/4):??, December 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332368.

Anonymous:1948:CMR

[Ano48c] Anonymous. Corrigenda: “Moments of the Ratio of the Mean Deviation to the Standard Deviation for Normal Samples”. Biometrika, 35(3/4):424, December 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332365. See [Gea36].

Anonymous:1948:CPM

[Ano48d] Anonymous. Corrigenda: “The Probability of the Mean Deviation”. Biometrika, 35(3/4):424, December 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332366. See [Gea36, Pea45].

Anonymous:1948:FMa

[Ano48e] Anonymous. Front matter. Biometrika, 35(1/2):??, May 1948. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332622.

Anonymous:1948:FMb

[Ano48f] Anonymous. Front matter. Biometrika, 35(3/4):??, December 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332341.

Anonymous:1948:VI

[Ano48g] Anonymous. Volume information. Biometrika, 35(3/4):i–vi, December 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332340.

Anonymous:1949:BMa

[Ano49a] Anonymous. Back matter. Biometrika, 36(1/2):??, June 1949. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332547.

Anonymous:1949:BMb

[Ano49b] Anonymous. Back matter. Biometrika, 36(3/4):??, December 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332688.

Anonymous:1949:FMa

[Ano49c] Anonymous. Front matter. Biometrika, 36(1/2):??, June 1949. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332523.

Anonymous:1949:FMb

[Ano49d] Anonymous. Front matter. Biometrika, 36(3/4):??, December 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332664.

Anonymous:1949:VI

[Ano49e] Anonymous. Volume information. Biometrika, 36(3/4):i–vi, December 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332663.

Anscombe:1948:TPB

[Ans48] F. J. Anscombe. The transformation of Poisson, binomial and negative-binomial data. Biometrika, 35(3/4):246–254, December 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332343.

Anscombe:1949:LST

[Ans49] F. J. Anscombe. Large-sample theory of sequential estimation. Biometrika, 36(3/4):455–458, December 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332681.

Armitage:1947:CSU

[Arm47] P. Armitage. A comparison of stratified with unrestricted random sampling from a finite population. Biometrika, 34(3/4):273–280, December 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332437.

Armitage:1949:OPA

[Arm49] P. Armitage. An overlap problem arising in particle counting. Biometrika, 36(3/4):257–266, December 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332666.

Aroian:1947:NCF

[Aro47] Leo A. Aroian. Note on the cumulants of Fisher’s z-distribution. Biometrika, 34(3/4):359–360, December 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332447.

Aspin:1948:EFD

[Asp48] Alice A. Aspin. An examination and further development of a formula arising in the problem of comparing two mean values. Biometrika, 35(1/2):88–96, May 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332631.

Aspin:1949:TUC

[AW49] Alice A. Aspin and B. L. Welch. Tables for use in comparisons whose accuracy involves two variances, separately estimated. Biometrika, 36(3/4):290–296, December 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332668.

Baldwin:1946:TPP

[Bal46] Elizabeth M. Baldwin. Table of percentage points of the tdistribution. Biometrika, 33(4):362, June 1946. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332199.

Bartlett:1940:NIQ

[Bar40] M. S. Bartlett. A note on the interpretation of quasi-sufficiency. Biometrika, 31(3/4):391–392, March 1940. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332618.

Bartlett:1941:SSC

[Bar41] M. S. Bartlett. The statistical significance of canonical correlations. Biometrika, 32(1):29–37, January 1941. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332247.

Barnard:1947:TNP

[Bar47a] G. A. Barnard. 2 × 2 tables. A note on E. S. Pearson’s paper. Biometrika, 34(1/2):168–169, January 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332519.

Barnard:1947:MSL

[Bar47b] G. A. Barnard. The meaning of a significance level. Biometrika, 34(1/2):179–182, January 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332521.

Barnard:1947:STT

[Bar47c] G. A. Barnard. Significance tests for 2 × 2 tables. Biometrika, 34(1/2):123–138, January 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332517.

Baten:1941:CAB

[Bat41] William Dowell Baten. A comparison of annual and biennial inflorescences of Daucus Carota (wild carrot). Biometrika, 32(1):82–87, January 1941. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332254.

Bateman:1948:PFL

[Bat48] G. Bateman. On the power function of the longest run as a test for randomness in a sequence of alternatives. Biometrika, 35(1/2): 97–112, May 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332632.

Bateman:1949:CFW

[Bat49] G. I. Bateman. The characteristic function of a weighted sum of non-central squares of normal variates subject to s linear restraints. Biometrika, 36(3/4):460–462, December 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332683.

Beall:1942:TDE

[Bea42] Geoffrey Beall. The transformation of data from entomological field experiments so that the analysis of variance becomes applicable. Biometrika, 32(3/4):243–262, April 1942. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332128.

Brownlee:1948:FRA

[BKL48] K. A. Brownlee, B. K. Kelly, and P. K. Loraine. Fractional replication arrangements for factorial experiments with factors at two levels. Biometrika, 35(3/4):268–276, December 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332346.

Brownlee:1948:RBF

[BL48] K. A. Brownlee and P. K. Loraine. The relationship between finite groups and completely orthogonal squares, cubes, and hyper-cubes.

Biometrika, 35(3/4):277–282, December 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332347.

Batrawi:1947:SFD

[BM47] A. Batrawi and G. M. Morant. A study of a First Dynasty series of Egyptian skulls from Sakkara and of an Eleventh Dynasty series from Thebes. Biometrika, 34(1/2):18–27, January 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332509.

Box:1949:GDT

[Box49] G. E. P. Box. A general distribution theory for a class of likelihood criteria. Biometrika, 36(3/4):317–346, December 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332671.

Carter:1947:APP

[Car47] A. H. Carter. Approximation to percentage points of the zdistribution. Biometrika, 34(3/4):352–358, December 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332446.

Carter:1949:ECR

[Car49] A. H. Carter. The estimation and comparison of residual regressions where there are two or more related sets of observations. Biometrika, 36(1/2):26–46, June 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332527.

Comrie:1941:TLC

[CH41] L. J. Comrie and H. O. Hartley. Table of Lagrangian coefficients for harmonic interpolation in certain tables of percentage points. Biometrika, 32(2):183–186, October 1941. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332209.

Chambers:1946:STA

[Cha46] E. G. Chambers. Statistical-techniques in applied psychology. Biometrika, 33(4):269–273, June 1946. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332191.

Cockayne:1940:TVO

[Coc40] E. A. Cockayne. Transposition of the viscera and other reversals of symmetry in monozygotic twins. Biometrika, 31(3/4):287–294, March 1940. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332610.

Cox:1948:NAD

[Cox48] D. R. Cox. A note on the asymptotic distribution of range. Biometrika, 35(3/4):310–315, December 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332353.

Daniels:1941:PDE

[Dan41] H. E. Daniels. A property of the distribution of extremes. Biometrika, 32(2):194–195, October 1941. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332213.

Daniels:1944:RBM

[Dan44] H. E. Daniels. The relation between measures of correlation in the universe of sample permutations. Biometrika, 33(2):129–135, August 1944. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334112.

Daniels:1948:PRC

[Dan48] H. E. Daniels. A property of rank correlations. Biometrika, 35 (3/4):416–417, December 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332360.

David:1940:BRB

[Dav40] F. N. David. Book review: The Principles of the Mathematical Theory of Correlation. Biometrika, 31(3/4):396, March 1940. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332620.

David:1947:STG

[Dav47a] F. N. David. A χ^{2} ‘smooth’ test for goodness of fit. Biometrika, 34(3/4):299–310, December 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332441.

David:1947:BRBa

[Dav47b] F. N. David. Book review: A First Course in Mathematical Statistics, by C. E. Weatherburn. Biometrika, 34(3/4):373, December 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332452.

David:1947:BRBb

[Dav47c] F. N. David. Book review: Mathematical Methods of Statistics, by H. Cramér. Biometrika, 34(3/4):374, December 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332454.

David:1947:PFT

[Dav47d] F. N. David. A power function for tests of randomness in a sequence of alternatives. Biometrika, 34(3/4):335–339, December 1947. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332444.

David:1948:CBC

[Dav48] F. N. David. Correlations between χ cells. Biometrika, 35 (3/4):418–422, December 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332362.

David:1949:BRBb

[Dav49a] F. N. David. Book review: The Fundamentals of Statistics, by T. L. Kelley. Biometrika, 36(3/4):470, December 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332687.

David:1949:BRBa

[Dav49b] F. N. David. Book review: Theory of Probability, by Harold Jeffreys. Biometrika, 36(1/2):236, June 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332545.

David:1949:MD

[Dav49c] F. N. David. The moments of the z and F distributions. Biometrika, 36(3/4):394–403, December 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332676.

David:1949:NAF

[Dav49d] F. N. David. Note on the application of Fisher’s k-statistics. Biometrika, 36(3/4):383–393, December 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332675.

David:1948:PIT

[DJ48] F. N. David and N. L. Johnson. The probability integral transformation when parameters are estimated from the sample. Biometrika, 35(1/2):182–190, May 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332638.

Daniels:1947:SRC

[DK47] H. E. Daniels and M. G. Kendall. The significance of rank correlations where parental correlation exists. Biometrika, 34(3/4):197–208, December 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332433.

David:1949:TSF

[DK49] F. N. David and M. G. Kendall. Tables of symmetric functions — Part I. Biometrika, 36(3/4):431–449, December 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332679.

Dumas:1948:CFK

[Dum48] M. Dumas. Sur les courbes de frequence de K. Pearson. (French) [On K. Pearson’s frequency curves]. Biometrika, 35(1/2):113–117, May 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332633.

Elfving:1947:ADR

[Elf47] G. Elfving. The asymptotical distribution of range in samples from a normal population. Biometrika, 34(1/2):111–119, January 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332515.

Finney:1947:EIR

[Fin47] D. J. Finney. The estimation from individual records of the relationship between dose and quantal response. Biometrika, 34 (3/4):320–334, December 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332443.

Finney:1948:FYT

[Fin48] D. J. Finney. The Fisher–Yates test of significance in 2 × 2 contingency tables. Biometrika, 35(1/2):145–156, May 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332635.

Finney:1949:EPT

[Fin49a] D. J. Finney. The estimation of the parameters of tolerance distributions. Biometrika, 36(3/4):239–256, December 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332665.

Finney:1949:MEF

[Fin49b] D. J. Finney. On a method of estimating frequencies. Biometrika, 36(1/2):233–234, June 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332543.

Fog:1948:GMT

[Fog48] David Fog. The geometrical method in the theory of sampling. Biometrika, 35(1/2):46–54, May 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332627.

Finney:1948:TCW

[FS48] D. J. Finney and W. L. Stevens. A table for the calculation of working probits and weights in probit analysis. Biometrika, 35(1/2):191– 201, May 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332639.

Garwood:1941:AML

[Gar41] F. Garwood. The application of maximum likelihood to dosagemortality curves. Biometrika, 32(1):46–58, January 1941. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332249.

Garwood:1947:VOG

[Gar47] F. Garwood. The variance of the overlap of geometrical figures with reference to a bombing problem. Biometrika, 34(1/2):1–17, January 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332508.

Gayen:1949:DSR

[Gay49] A. K. Gayen. The distribution of ‘Student’s t in random samples of any size drawn from non-normal universes. Biometrika, 36 (3/4):353–369, December 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332673.

Geary:1936:MRM

[Gea36] R. C. Geary. Moments of the ratio of the mean deviation to the standard deviation for normal samples. Biometrika, 28(3/4):295–307, December 1936. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333953.

See corrigenda [Ano48c] and comments [SBL79].

Geary:1943:MRQ

[Gea43] R. C. Geary. Minimum range for quasi-normal distributions. Biometrika, 33(1):100–103, April 1943. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333623.

Geary:1944:CCE

[Gea44] R. C. Geary. Comparison of the concepts of efficiency and closeness for consistent estimates of a parameter. Biometrika, 33(2):123–128, August 1944. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334111.

Geary:1947:FDS

[Gea47a] R. C. Geary. The frequency distribution of \sqrt{b_{1}} for samples of all sizes drawn a random from a normal population. Biometrika, 34(1/2):68– 97, January 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332513.

Geary:1947:TN

[Gea47b] R. C. Geary. Testing for normality. Biometrika, 34(3/4):209–242, December 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332434.

Ghurye:1949:UST

[Ghu49] S. G. Ghurye. On the use of Student’s t-test in an asymmetrical population. Biometrika, 36(3/4):426–430, December 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332678.

Goodman:1940:HRI

[GM40] C. N. Goodman and G. M. Morant. The human remains of the Iron Age and other periods from Maiden Castle, Dorset. Biometrika, 31(3/4):295–312, March 1940. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332611.

Godwin:1945:DEM

[God45] H. J. Godwin. On the distribution of the estimate of mean deviation obtained from samples from a normal population. Biometrika, 33(3):254–256, November 1945. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332305.

Godwin:1948:FNM

[God48] H. J. Godwin. A further note on the mean deviation. Biometrika, 35(3/4):304–309, December 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332352.

Godwin:1949:EDL

[God49] H. J. Godwin. On the estimation of dispersion by linear systematic statistics. Biometrika, 36(1/2):92–100, June 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332533.

Greenwood:1941:MSG

[Gre41] Major Greenwood. Medical statistics from Graunt to Farr. Biometrika, 32(2):101–127, October 1941. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332206.

Greenwood:1942:MSG

[Gre42] Major Greenwood. Medical statistics from Graunt to Farr. Biometrika, 32(3/4):203–225, April 1942. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332126.

Greenwood:1943:MSG

[Gre43] Major Greenwood. Medical statistics from Graunt to Farr (concluded). Biometrika, 33(1):1–24, April 1943. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333616.

Greenwood:1948:JFT

[Gre48] Major Greenwood. James Fowler Tocher. Biometrika, 35(1/2):1–5, May 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332623.

Greenwood:1949:IM

[Gre49] Major Greenwood. The infectiousness of measles. Biometrika, 36(1/2):1–8, June 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332524.

Gumbel:1942:STG

[Gum42] E. J. Gumbel. Simple tests for given hypotheses. Biometrika, 32(3/4):317–333, April 1942. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332136.

Gumbel:1949:PTR

[Gum49] E. J. Gumbel. Probability tables for the range. Biometrika, 36(1/2):142–148, June 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332538.

Geary:1947:CUM

[GW47] R. C. Geary and J. P. G. Worlledge. On the computation of universal moments of tests of statistical normality derived from samples drawn at random from a normal universe. Application to the calculation of the seventh moment of b_{2}. Biometrika, 34(1/2):98–110, January 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332514.

See corrigenda [GW50].

Geary:1950:CCU

[GW50] R. C. Geary and J. P. G. Worlledge. Corrigenda: “On the Computation of Universal Moments of Tests of Statistical Normality Derived from Samples Drawn at Random from a Normal Universe. Application to the Calculation of the Seventh Moment of b_{2}”. Biometrika, 37(1/2):189, June 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332167. See [GW47].

H:1947:BRB

[H.47] J. B. S. H. Book review: Advances in Genetics. Biometrika, 34 (3/4):373–374, December 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332453.

Haldane:1940:CMB

[Hal40a] J. B. S. Haldane. The cumulants and moments of the binomial distribution, and the cumulants of χ^{2} for a (n × 2)-fold table. Biometrika, 31(3/4):392–396, March 1940. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332619.

Haldane:1940:MVW

[Hal40b] J. B. S. Haldane. The mean and variance of χ^{2}, when used as a test of homogeneity, when expectations are small. Biometrika, 31(3/4):346– 355, March 1940. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332614.

Haldane:1941:CDS

[Hal41] J. B. S. Haldane. The cumulants of the distribution of the square of a variate. Biometrika, 32(2):199–200, October 1941. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332215.

Haldane:1942:MMN

[Hal42a] J. B. S. Haldane. The mode and median of a nearly normal distribution with given cumulants. Biometrika, 32(3/4):294–299, April 1942. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332132.

Haldane:1942:MDP

[Hal42b] J. B. S. Haldane. Moments of the distributions of powers and products of normal variates. Biometrika, 32(3/4):226–242, April 1942. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332127.

Haldane:1945:MFT

[Hal45a] J. B. S. Haldane. Moments of r and χ^{2} for a fourfold table in the absence of association. Biometrika, 33(3):231–233, November 1945. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332301.

Haldane:1945:MEF

[Hal45b] J. B. S. Haldane. On a method of estimating frequencies. Biometrika, 33(3):222–225, November 1945. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332299.

Haldane:1945:UTH

[Hal45c] J. B. S. Haldane. The use of χ^{2} as a test of homogeneity in a (n×2)- fold table when expectations are small. Biometrika, 33(3):234–238, November 1945. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332302.

Haldane:1948:NMM

[Hal48a] J. B. S. Haldane. Note on the median of a multivariate distribution. Biometrika, 35(3/4):414–415, December 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332359.

Haldane:1948:POV

[Hal48b] J. B. S. Haldane. The precision of observed values of small frequencies. Biometrika, 35(3/4):297–300, December 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332350.

Haldane:1949:NNN

[Hal49] J. B. S. Haldane. A note on non-normal correlation. Biometrika, 36(3/4):467–468, December 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332685.

Hartley:1940:THS

[Har40] H. O. Hartley. Testing the homogeneity of a set of variances. Biometrika, 31(3/4):249–255, March 1940. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332607.

Hartley:1942:RRS

[Har42] H. O. Hartley. The range in random samples. Biometrika, 32(3/4):334–348, April 1942. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332137.

Hartley:1944:SES

[Har44] H. O. Hartley. Studentization or the elimination of the standard deviation of the parent population from the random sample-distribution of statistics. Biometrika, 33(2):173–180, August 1944. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334115.

Hartley:1945:ANC

[Har45] H. O. Hartley. Appendix: Note on the calculation of the distribution of the estimate of mean deviation in normal samples. Biometrika, 33(3):257–258, November 1945. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332306.

Hartley:1948:AED

[Har48a] H. O. Hartley. Approximation errors in distributions of independent variates. Biometrika, 35(3/4):417–418, December 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332361.

Hartley:1948:ENL

[Har48b] H. O. Hartley. The estimation of non-linear parameters by ‘internal least squares’. Biometrika, 35(1/2):32–45, May 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332626.

Hartley:1949:TSH

[Har49] H. O. Hartley. Tests of significance in harmonic analysis. Biometrika, 36(1/2):194–201, June 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332541.

Hartley:1947:NSP

[HK47] H. O. Hartley and S. H. Khamis. A numerical solution of the problem of moments. Biometrika, 34(3/4):340–351, December 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332445.

Hsu:1940:DFS

[HL40] C. T. Hsu and D. N. Lawley. The derivation of the fifth and sixth moments of the distribution of b_{2} in samples from a normal population. Biometrika, 31(3/4):238–248, March 1940. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332606.

Hoffding:1947:DRC

[Hof47] Wassily Hoffding. On the distribution of the rank correlation coefficient τ when the variates are not independent. Biometrika, 34 (3/4):183–196, December 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332432.

Hacker:1944:GSW

[HP44] H. P. Hacker and H. S. Pearson. The growth, survival, wandering and variation of the long-tailed field mouse, Apodemus Sylvaticus. Biometrika, 33(2):136–162, August 1944. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334113.

Hacker:1946:GSW

[HP46] H. P. Hacker and H. S. Pearson. The growth, survival, wandering and variation of the long-tailed field mouse, Apodemus Sylvaticus. Biometrika, 33(4):333–361, June 1946. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332198.

Hsu:1940:GAV

[Hsu40] P. L. Hsu. On generalized analysis of variance. (I). Biometrika, 31(3/4):221–237, March 1940. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332605.

Hsu:1941:AVP

[Hsu41a] P. L. Hsu. Analysis of variance from the power function standpoint. Biometrika, 32(1):62–69, January 1941. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332251.

Hsu:1941:LDC

[Hsu41b] P. L. Hsu. On the limiting distribution of the canonical correlations. Biometrika, 32(1):38–45, January 1941. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332248.

Huntington:1940:SFR

[Hun40] E. V. Huntington. Stirling’s formula with remainder. Biometrika, 31(3/4):390, March 1940. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332617.

Huzurbazar:1949:PDA

[Huz49] V. S. Huzurbazar. On a property of distributions admitting sufficient statistics. Biometrika, 36(1/2):71–74, June 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332531.

Irwin:1949:NSC

[Irw49] J. O. Irwin. A note on the subdividsion of χ^{2} into components. Biometrika, 36(1/2):130–134, June 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332536.

Iyer:1949:FSM

[Iye49] P. V. Krishna Iyer. The first and second moments of some probability distributions arising from points on a lattice and their application. Biometrika, 36(1/2):135–141, June 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332537.

J:1948:BRB

[J.48] N. L. J. Book review: Probit Analysis, by D. J. Finney. Biometrika, 35(3/4):423, December 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332364.

Johnson:1948:ASA

[Joh48a] N. L. Johnson. Alternative systems in the analysis of variance. Biometrika, 35(1/2):80–87, May 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332630.

Johnson:1948:TSV

[Joh48b] N. L. Johnson. Tests of significance in the variate difference method. Biometrika, 35(1/2):206–209, May 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332642.

Johnson:1949:BDB

[Joh49a] N. L. Johnson. Bivariate distributions based on simple translation systems. Biometrika, 36(3/4):297–304, December 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332669.

Johnson:1949:SFC

[Joh49b] N. L. Johnson. Systems of frequency curves generated by methods of translation. Biometrika, 36(1/2):149–176, June 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332539.

Jones:1946:UMR

[Jon46] A. E. Jones. A useful method for the routine estimation of dispersion from large samples. Biometrika, 33(4):274–282, June 1946. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332192.

Jones:1948:SSC

[Jon48] A. E. Jones. Systematic sampling of continuous parameter populations. Biometrika, 35(3/4):283–290, December 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332348.

Johnson:1940:ANC

[JW40] N. L. Johnson and B. L. Welch. Applications of the non-central t-distribution. Biometrika, 31(3/4):362–389, March 1940. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332616.

Kempthorne:1947:SAC

[Kem47] O. Kempthorne. A simple approach to confounding and fractional replication in factorial experiments. Biometrika, 34(3/4):255–272, December 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332436.

Kendall:1941:PRC

[Ken41a] M. G. Kendall. Proof of relations connected with the tetrachoric series and its generalization. Biometrika, 32(2):196–198, October 1941. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332214.

Kendall:1941:RRS

[Ken41b] M. G. Kendall. A recurrence relation for the semi-invariants of Pearson curves. Biometrika, 32(1):81–82, January 1941. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332253.

Kendall:1941:TR

[Ken41c] M. G. Kendall. A theory of randomness. Biometrika, 32(1):1–15, January 1941. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332245.

Kendall:1942:PRC

[Ken42] M. G. Kendall. Partial rank correlation. Biometrika, 32(3/4):277– 283, April 1942. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332130.

Kendall:1944:ATS

[Ken44] M. G. Kendall. On autoregressive time series. Biometrika, 33(2):105– 122, August 1944. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334110.

Kendall:1945:TTR

[Ken45] M. G. Kendall. The treatment of ties in ranking problems. Biometrika, 33(3):239–251, November 1945. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332303.

Kendall:1947:VWB

[Ken47] M. G. Kendall. The variance of τ when both rankings contain ties. Biometrika, 34(3/4):297–298, December 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332440.

Kendall:1948:SMP

[Ken48a] David G. Kendall. On some modes of population growth leading to R. A. Fisher’s logarithmic series distribution. Biometrika, 35(1/2): 6–15, May 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332624.

Kendall:1948:RVG

[Ken48b] David G. Kendall. On the role of variable generation time in the development of a stochastic birth process. Biometrika, 35 (3/4):316–330, December 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332354.

Kendall:1948:CDJ

[Ken48c] M. G. Kendall. Continuation of Dr Jones’s paper. Biometrika, 35 (3/4):291–296, December 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332349.

Kendall:1949:RTP

[Ken49a] M. G. Kendall. On the reconciliation of theories of probability. Biometrika, 36(1/2):101–116, June 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332534.

Kendall:1949:RPM

[Ken49b] M. G. Kendall. Rank and product-moment correlation. Biometrika, 36(1/2):177–193, June 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332540.

Kendall:1949:TAS

[Ken49c] M. G. Kendall. Tables of autoregressive series. Biometrika, 36 (3/4):267–289, December 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332667.

Kendall:1940:MPC

[KS40] M. G. Kendall and B. Babington Smith. On the method of paired comparisons. Biometrika, 31(3/4):324–345, March 1940. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332613.

Lancaster:1949:CPA

[Lan49a] H. O. Lancaster. The combination of probabilities arising from data in discrete distributions. Biometrika, 36(3/4):370–382, December 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332674. See corri-

genda [Lan50a].

Lancaster:1949:DPC

[Lan49b] H. O. Lancaster. The derivation and partition of χ^{2} in certain discrete distributions. Biometrika, 36(1/2):117–129, June 1949. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332535. See corrigenda

[Lan50b].

Lancaster:1950:CCP

[Lan50a] H. O. Lancaster. Corrigenda: “The Combination of Probabilities Arising from Data in Discrete Distributions”. Biometrika, 37(3/4):452, December 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332400. See [Lan49a].

Lancaster:1950:CDP

[Lan50b] H. O. Lancaster. Corrigenda: “The Derivation and Partition of χ^{2} in Certain Discrete Distributions”. Biometrika, 37(3/4):452, December 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332399. See

[Lan49b].

Leser:1942:IMF

[Les42] C. E. V. Leser. Inequalities for multivariate frequency distributions. Biometrika, 32(3/4):284–293, April 1942. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332131.

Leslie:1945:UMC

[Les45] P. H. Leslie. On the use of matrices in certain population mathematics. Biometrika, 33(3):183–212, November 1945. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332297.

Leslie:1948:SFN

[Les48] P. H. Leslie. Some further notes on the use of matrices in population mathematics. Biometrika, 35(3/4):213–245, December 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332342.

Lindley:1946:SSE

[Lin46] D. V. Lindley. On the solution of some equations in least squares. Biometrika, 33(4):326–327, June 1946. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332196.

Little:1943:SSH

[Lit43] K. L. Little. A study of a series of human skulls from Castle Hill, Scarborough. Biometrika, 33(1):25–35, April 1943. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333617.

Lord:1947:URP

[Lor47] E. Lord. The use of range in place of standard deviation in the t-test. Biometrika, 34(1/2):41–67, January 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor. org/stable/2332512. See corrigenda [Lor52].

Lord:1952:CUR

[Lor52] E. Lord. Corrigenda: “The Use of Range in Place of Standard Deviation in the t-Test”. Biometrika, 39(3/4):442, December 1952. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334045. See [Lor47].

M:1941:BRB

[M.41] G. M. M. Book review: A Bibliography of Human Morphology, 1914– 1939, by Wilton M. Krogman. Biometrika, 32(2):194, October 1941. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332212.

Merrington:1941:NAP

[Mer41] Maxine Merrington. Numerical approximations to the percentage points of the χ^{2} distribution. Biometrika, 32(2):200–202, October 1941. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332216.

Merrington:1942:TPP

[Mer42] Maxine Merrington. Table of percentage points of the t-distribution. Biometrika, 32(3/4):300, April 1942. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332133.

Moore:1949:TRS

[Moo49] P. G. Moore. A test for randomness in a sequence of two alternatives involving a 2×2 table. Biometrika, 36(3/4):305–316, December 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332670.

Moran:1947:MPC

[Mor47a] P. A. P. Moran. On the method of paired comparisons. Biometrika, 34(3/4):363–365, December 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332449.

Moran:1947:STT

[Mor47b] P. A. P. Moran. Some theorems on time series. I. Biometrika, 34 (3/4):281–291, December 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332438.

Moran:1948:RCP

[Mor48a] P. A. P. Moran. Rank correlation and product-moment correlation. Biometrika, 35(1/2):203–206, May 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332641.

Moran:1948:STT

[Mor48b] P. A. P. Moran. Some theorems on time series: II the significance of the serial correlation coefficient. Biometrika, 35(3/4):255–260, December 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332344.

Morant:1948:CHW

[Mor48c] G. M. Morant. Comparisons of heights and weights of German civilians recorded in 1946–7 and Royal Air Force and other British series. Biometrika, 35(3/4):368–396, December 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332357.

Moran:1949:STD

[Mor49] P. A. P. Moran. The spectral theory of discrete stochastic processes. Biometrika, 36(1/2):63–70, June 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332530.

Merrington:1943:TPP

[MT43] Maxine Merrington and Catherine M. Thompson. Tables of percentage points of the inverted beta (f) distribution. Biometrika, 33(1): 73–88, April 1943. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333621.

Nair:1947:NMD

[Nai47] K. R. Nair. A note on the mean deviation from the median. Biometrika, 34(3/4):360–362, December 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332448.

Nair:1948:DED

[Nai48a] K. R. Nair. The distribution of the extreme deviate from the sample mean and its Studentized form. Biometrika, 35(1/2):118–144, May 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332634.

Nair:1948:SFE

[Nai48b] K. R. Nair. The Studentized form of the extreme mean square test in the analysis of variance. Biometrika, 35(1/2):16–31, May 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332625.

Nair:1949:FNM

[Nai49] K. R. Nair. A further note on the mean deviation from the median. Biometrika, 36(1/2):234–235, June 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332544.

Neyman:1941:FAT

[Ney41] J. Neyman. Fiducial argument and the theory of confidence intervals. Biometrika, 32(2):128–150, October 1941. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332207.

Nicholson:1941:GAF

[Nic41] C. Nicholson. A geometrical analysis of the frequency distribution of the ratio between two variables. Biometrika, 32(1):16–28, January 1941. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332246.

Nicholson:1943:PIT

[Nic43] C. Nicholson. The probability integral for two variables. Biometrika, 33(1):59–72, April 1943. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333620.

Orcutt:1948:TSC

[OJ48] G. H. Orcutt and S. F. James. Testing the significance of correlation between time series. Biometrika, 35(3/4):397–413, December 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332358.

P:1942:BRB

[P.42] E. S. P. Book review: The Second Yearbook of Research and Statistical Methodology Books and Reviews by Oscar Krisen Buros. Biometrika, 32(3/4):349–350, April 1942. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332138.

Patnaik:1948:PFT

[Pat48] P. B. Patnaik. The power function of the test for the difference between two proportions in a 2 × 2 table. Biometrika, 35(1/2):157– 175, May 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332636.

See corrigenda [Pat59].

Patnaik:1949:NCD

[Pat49] P. B. Patnaik. The non-central χ^{2}- and F -distribution and their applications. Biometrika, 36(1/2):202–232, June 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332542.

Patnaik:1959:CPF

[Pat59] P. B. Patnaik. Corrigenda: “The Power Function of the Test for the Difference Between Two Proportions in a 2 × 2 Table”. Biometrika, 46(3/4):502, December 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333568. See [Pat48].

Plackett:1946:DOM

[PB46] R. L. Plackett and J. P. Burman. The design of optimum multifactorial experiments. Biometrika, 33(4):305–325, June 1946. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332195.

Pearson:1941:NFP

[Pea41] E. S. Pearson. A note on further properties of statistical tests. Biometrika, 32(1):59–61, January 1941. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332250.

Pearson:1942:NTS

[Pea42] E. S. Pearson. Notes on testing statistical hypotheses. Biometrika, 32(3/4):311–316, April 1942. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332135.

Pearson:1945:EN

[Pea45] E. S. Pearson. Editorial note. Biometrika, 33(3):252–253, November 1945. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332304. See

corrigenda [Ano48d].

Pearson:1947:CST

[Pea47] E. S. Pearson. The choice of statistical tests illustrated on the interpretation of data classed in a 2 × 2 table. Biometrika, 34 (1/2):139–167, January 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332518.

Pearson:1948:NPH

[Pea48] E. S. Pearson. Note on Professor Haldane’s paper regarding the treatment of rare events. Biometrika, 35(3/4):301–303, December 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332351.

Pearson:1942:PIR

[PH42] E. S. Pearson and H. O. Hartley. The probability integral of the range in samples of n observations from a normal population. Biometrika, 32(3/4):301–310, April 1942. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332134.

Pearson:1943:TPI

[PH43] E. S. Pearson and H. O. Hartley. Tables of the probability integral of the Studentized range. Biometrika, 33(1):89–99, April 1943. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333622.

Plackett:1946:SGM

[Pla46] R. L. Plackett. Some generalizations in the multifactorial design. Biometrika, 33(4):328–332, June 1946. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332197.

Plackett:1947:ETE

[Pla47a] R. L. Plackett. An exact test for the equality of variances. Biometrika, 34(3/4):311–319, December 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332442.

Plackett:1947:LRM

[Pla47b] R. L. Plackett. Limits of the ratio of mean range to standard deviation. Biometrika, 34(1/2):120–122, January 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332516.

Plackett:1949:HNM

[Pla49] R. L. Plackett. A historical note on the method of least squares. Biometrika, 36(3/4):458–460, December 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332682.

Pearson:1948:TPF

[PM48] E. S. Pearson and Maxine Merrington. 2 × 2 tables; the power function of the test on a randomized experiment. Biometrika, 35 (3/4):331–345, December 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332355.

Przyborowski:1940:HRT

[PW40] J. Przyborowski and H. Wilenski. Homogeneity of results in testing samples from Poisson series: With an application to testing clover seed for dodder. Biometrika, 31(3/4):313–323, March 1940. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332612.

Quenouille:1947:NCA

[Que47] M. H. Quenouille. Notes on the calculation of autocorrelations of linear autoregressive schemes. Biometrika, 34(3/4):365–367, December 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332450.

Quenouille:1948:SRT

[Que48] M. H. Quenouille. Some results in the testing of serial correlation coefficients. Biometrika, 35(3/4):261–267, December 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332345.

Quenouille:1949:MTE

[Que49] M. H. Quenouille. On a method of trend elimination. Biometrika, 36(1/2):75–91, June 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332532.

Rao:1948:TSM

[Rao48] C. Radhakrishna Rao. Tests of significance in multivariate analysis. Biometrika, 35(1/2):58–79, May 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332629.

Rendel:1943:VWH

[Ren43] J. M. Rendel. Variations in the weights of hatched and unhatched ducks’ eggs. Biometrika, 33(1):48–58, April 1943. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333619.

Russel:1945:ERD

[Rus45] C. Scott Russel. Errors in the routine daily measurement of the puerperal uterus. Biometrika, 33(3):213–221, November 1945. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332298.

Shenton:1979:CPR

[SBL79] L. R. Shenton, K. O. Bowman, and H. K. Lam. Comments on a paper by R. C. Geary on standardized mean deviation. Biometrika, 66(2):400–401, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335679. See [Gea36].

Schreiner:1940:CPD

[Sch40] Alette Schreiner. Certain projective depth and breadth measurements of the facial skeleton in man. Biometrika, 31(3/4):272–286, March 1940. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332609.

Scheffe:1944:NUT

[Sch44] Henry Scheffé. Note on the use of the tables of percentage points of the incomplete beta function to calculate small sample confidence intervals for a binomial p. Biometrika, 33(2):181, August 1944. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334116.

Seal:1945:MPC

[Sea45] H. L. Seal. The mathematics of a population composed of k stationary strata each recruited from the stratum below and supported at the lowest level by a uniform annual number of entrants. Biometrika, 33(3):226–230, November 1945. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332300.

Seal:1948:NST

[Sea48] H. L. Seal. A note on the χ^{2} smooth test. Biometrika, 35(1/2):202, May 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332640.

Shenton:1949:EMM

[She49] L. R. Shenton. On the efficiency of the method of moments and Neyman’s type a distribution. Biometrika, 36(3/4):450–454, December 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332680.

Sichel:1949:MFM

[Sic49] Herbert S. Sichel. The method of frequency-moments and its application to Type VII populations. Biometrika, 36(3/4):404–425, December 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332677.

Sillitto:1947:DKC

[Sil47] G. P. Sillitto. The distribution of Kendall’s τ coefficient of rank correlation in rankings containing tie. Biometrika, 34(1/2):36–40, January 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332511.

Sillitto:1949:NAP

[Sil49] G. P. Sillitto. Note on approximations to the power function of the 2 × 2 comparative trial’. Biometrika, 36(3/4):347–352, December 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332672.

Simaika:1941:OPT

[Sim41] J. B. Simaika. On an optimum property of two important statistical tests. Biometrika, 32(1):70–80, January 1941. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332252.

Simaika:1942:IFP

[Sim42] J. B. Simaika. Interpolation for fresh probability levels between the standard table levels of a function. Biometrika, 32(3/4):263–276, April 1942. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332129.

Smith:1949:BRB

[Smi49] Cedric A. B. Smith. Book review: Probability Thoery for Statistical Methods, by F. N. David. Biometrika, 36(3/4):469–470, December 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332686.

Stevens:1948:SAN

[Ste48] W. L. Stevens. Statistical analysis of a non-orthogonal tri-factorial experiment. Biometrika, 35(3/4):346–367, December 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332356.

Thompson:1941:TPPb

[Tho41] Catherine M. Thompson. Table of percentage points of the χ^{2} distribution. Biometrika, 32(2):187–191, October 1941. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332210.

Thomson:1947:AFP

[Tho47] D. Halton Thomson. Approximate formulae for the percentage points of the incomplete beta function and of the χ^{2} distribution. Biometrika, 34(3/4):368–372, December 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332451.

Thomas:1949:GPB

[Tho49] Marjorie Thomas. A generalization of Poisson’s binomial limit for use in ecology. Biometrika, 36(1/2):18–25, June 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332526.

Tippett:1944:CIP

[Tip44] L. H. C. Tippett. The control of industrial processes subject to trends in quality. Biometrika, 33(2):163–172, August 1944. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334114.

Thompson:1946:TTH

[TM46] Catherine M. Thompson and Maxine Merrington. Tables for testing the homogeneity of a set of estimated variances. Biometrika, 33(4):296–304, June 1946. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2332194. With prefatory note by H. O. Hartley and E. S. Pearson.

Tocher:1949:NAG

[Toc49] K. D. Tocher. A note on the analysis of grouped probit data. Biometrika, 36(1/2):9–17, June 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332525.

Thompson:1941:TPPa

[TPCH41] Catherine M. Thompson, E. S. Pearson, L. J. Comrie, and H. O. Hartley. Tables of percentage points of the incomplete beta-function. Biometrika, 32(2):151–181, October 1941. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332208.

Truksa:1940:SDS

[Tru40] L. Truksa. The simultaneous distribution in samples of mean and standard deviation, and of mean and variance. Biometrika, 31(3/4):256–271, March 1940. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332608.

Vajda:1947:SSS

[Vaj47] S. Vajda. The stratified semi-stationary population. Biometrika, 34(3/4):243–254, December 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332435.

W:1948:BRB

[W.48] B. L. W. Book review: The Advanced Theory of Statistics by M. G. Kendall. Biometrika, 35(1/2):210–211, May 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332643.

Welch:1945:BRB

[Wel45] B. L. Welch. Book review: The Advanced Theory of Statistics, by Maurice G. Kendall. Biometrika, 33(3):266–267, November 1945. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332308.

Welch:1947:GSP

[Wel47] B. L. Welch. The generalization of ‘Student’s’ problem when several different population variances are involved. Biometrika, 34(1/2):28– 35, January 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332510.

Whitfield:1947:RCB

[Whi47] J. W. Whitfield. Rank correlation between two variables, one of which is ranked, the other dichotomous. Biometrika, 34(3/4): 292–296, December 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332439.

Whitfield:1949:ICR

[Whi49] J. W. Whitfield. Intra-class rank correlation. Biometrika, 36 (3/4):463–467, December 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332684.

Williams:1940:NSA

[Wil40] C. B. Williams. A note on the statistical analysis of sentence-length as a criterion of literary style. Biometrika, 31(3/4):356–361, March 1940. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332615.

Wilks:1941:BRB

[Wil41] S. S. Wilks. Book review: Theory of Probability by Harold Jeffreys. Biometrika, 32(2):192–194, October 1941. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332211.

Winsten:1946:ITM

[Win46] C. B. Winsten. Inequalities in terms of mean range. Biometrika, 33(4):283–295, June 1946. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332193.

Wishart:1947:CLD

[Wis47] John Wishart. The cumulants of the Z and of the logarithmic χ^{2} and t distributions. Biometrika, 34(1/2):170–178, January 1947. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332520. See corrigenda [Ano47c].

Wishart:1948:NPD

[Wis48a] John Wishart. Note on ‘proofs of the distribution law of the second order moment statistics’. Biometrika, 35(3/4):422, December 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332363.

Wishart:1948:PDL

[Wis48b] John Wishart. Proofs of the distribution law of the second order moment statistics. Biometrika, 35(1/2):55–57, May 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332628.

Wishart:1949:CMM

[Wis49] John Wishart. Cumulants of multivariate multinomial distributions. Biometrika, 36(1/2):47–58, June 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332528.

Woo:1943:SCH

[Woo43] T. L. Woo. A study of the Chinese humerus. Biometrika, 33(1):36– 47, April 1943. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333618.

Yates:1948:ACT

[Yat48a] F. Yates. The analysis of contingency tables with groupings based on quantitative characters. Biometrika, 35(1/2):176–181, May 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332637. See corri-

genda [Yat48b].

Yates:1948:CAC

[Yat48b] F. Yates. Corrigenda: “The Analysis of Contingency Tables with Groupings Based on Quantitative Characters”. Biometrika, 35(3/4):424, December 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332367. See [Yat48a].

Yule:1949:BRB

[Yul49] G. Udny Yule. Book review: Karl Pearson’s Early Statistical Papers, by Karl Pearson. Biometrika, 36(1/2):236–238, June 1949. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332546.